# Skip-Vision: Efficient and Scalable Acceleration of Vision-Language Models via Adaptive Token Skipping

### v2

@@ 1 Introduction
We systematically examine the parameter count and FLOPs distribution across different components of the LLM. As shown in Table 1, using LLaMA-3 [28] as an example, the parameter and FLOPS requirements for the FFN component are significantly higher than those for the attention module. Furthermore, as illustrated in Figure 2, when the number of visual tokens is large, the majority of computations are concentrated on FFN of these tokens. In particular, we computed the difference in the magnitudes of the features before and after the FFN transformation for different tokens. As illustrated in the figure 3, we observe that the updates introduced by the FFN to the visual tokens are significantly smaller than those that affect the text tokens. This observation suggests a potential approach: could we bypass FFN computations for a substantial portion of visual tokens? Furthermore, in the autoregressive training framework of MLLMs, only the final visual token utilizes its last hidden state for next-token prediction, while the other visual tokens primarily serve to convey causal information. During model inference, inspired by [7], we observe that most visual information is progressively integrated into the final token through the early causal layers. This raises the possibility of reducing reliance on the full visual scaling KV cache throughout inference, potentially accelerating the inference process.
@@
In summary, our contributions are as follows.
  • We introduce a novel and efficient Skip-Vision framework, using token merge and skip FFN strategy during training to reduce redundant computation for visual tokens.
@@ 2.2 Efficient multimodal large language models
To enhance the efficiency of vision-language models, one of the primary and widely employed approaches in contemporary research involves reducing the size of the backbone models [26, 123, 130, 36, 133, 83]. Simultaneously, given that visual tokens contribute significantly to computational demands, research focused on compressing or refining these tokens has garnered increasing attention. Such methods are often coupled with the design of vision-language projectors. For instance, techniques like the Perceiver Resampler [41, 6, 3] and Q-Former [58] utilize transformer-based mechanisms to consolidate visual tokens into more compact query sets. In this way, the transformer outputs corresponding to the positions of the learnable latent queries serve as the aggregated representation of the visual features [42]. Honeybee [14] propose two visual projectors, namely C-Abstractor and D-Abstractor. LLaVA-PruMerge [92] and MADTP [12] introduce adaptive methods for reducing visual tokens. In addition, while the large language model (LLM) field has employed various techniques for token reduction to accelerate inference and compress KV cache [34, 128], research in this area remains limited within the MLLM domain.
@@ 4.3 Inference speed-up: skip KV-cache
As depicted in Figures 4 b) and c), given the autoregressive nature of LLMs, earlier token information is progressively transferred to later tokens. Skip FFN further enhances this integration by suppressing the attention weights of skipped tokens, forcing the model to focus on key tokens and improving information aggregation efficiency. Empirical results show that, due to the inclusion of summary tokens, this omission does not compromise model performance, underscoring a distinctive advantage of our framework for more efficient inference.
@@ 5 Theoretical analysis of Skip-Vision
Bounded Lipschitz. If $W_{Q}$, $W_{K}$, $W_{V}$, $W_{1}$, $W_{2}$ are orthogonal matrices (spectral norm = 1), then:
@@ 6.4 Training setup
For LLaVA and LLaVA-HD setting, we use input resolutions of 336 and a dynamic resolution ranging from 336 to 672, respectively, with CLIP ViT-L/336px [87] as the vision encoder and LLaMA3 8B Instruct [28] as the backbone. For CoS setting, pretraining is conducted at 224 resolution, training only the multi-scale visual resampler with CLIP ViT-L/224px [87]. The model processes 80 visual tokens, including 16 global and 64 local tokens, and scales to 336 and 1296 tokens during SFT. In SFT, the resolution increases to 448, and all model parameters are unfrozen for full optimization.

### v3

@@ Abstract
00footnotetext: *This work was done when Weili Zeng interned at Ant Group.00footnotetext: Corresponding author
@@ 1 Introduction
We systematically examine the parameter count and FLOPs distribution across different components of the LLM. As shown in Table 1, using LLaMA-3 [29] as an example, the parameter and FLOPS requirements for the FFN are significantly higher than the attention module. Furthermore, as illustrated in Figure 2, when the number of visual tokens is large, the majority of computations are concentrated on FFN of these tokens. In particular, we computed the difference in the magnitudes of the features before and after the FFN transformation for different tokens. As illustrated in the figure 3, we observe that the updates introduced by the FFN to the visual tokens are significantly smaller than those that affect the text tokens. This is similar to the observation in [124]. This observation suggests a potential approach: could we bypass FFN computations for a substantial portion of visual tokens? Furthermore, in the autoregressive training framework of MLLMs, only the final visual token utilizes its last hidden state for next-token prediction, while the other visual tokens primarily serve to convey causal information. During model inference, inspired by [7], we observe that most visual information is progressively integrated into the final token through the early causal layers. This raises the possibility of reducing reliance on the full visual scaling KV cache throughout inference, potentially accelerating the inference process.
@@
In summary, Skip-Vision is designed to be seamlessly integrated into the standard SFT pipeline of MLLMs without introducing additional retraining or decoupled modules. It directly modifies the transformer’s computation flow, offering a practical and theoretically grounded acceleration solution for MLLM training and inference jointly. Our contributions are as follows:
  • We introduce a novel and efficient framework, using token merge and skip FFN strategy during training to reduce redundant computation for visual tokens.
@@ 2.2 Efficient multimodal large language models
To enhance the efficiency of vision-language models, one of the primary and widely employed approaches in contemporary research involves reducing the size of the backbone models [27, 125, 134, 37]. Simultaneously, given that visual tokens contribute significantly to computational demands, research focused on compressing or refining these tokens has garnered increasing attention. Such methods are often coupled with the design of vision-language projectors. For instance, techniques like the Perceiver Resampler [42, 6, 3] and Q-Former [59] utilize transformer-based mechanisms to consolidate visual tokens into more compact query sets. In this way, the transformer outputs corresponding to the positions of the learnable latent queries serve as the aggregated representation of the visual features [43]. Honeybee [14] propose two visual projectors, namely C-Abstractor and D-Abstractor. LLaVA-PruMerge [92] and MADTP [12] introduce adaptive methods for reducing visual tokens. In addition, while the large language model (LLM) field has employed various techniques for token reduction to accelerate inference and compress KV cache [35, 132], research in this area remains limited within the MLLM domain.
@@ 4.3 Inference speed-up: skip KV-cache
Equation 7 indicates that visual tokens skipped by FFN are removed from the KV cache after the pre-filling stage. As depicted in Figures 4 b) and c), given the autoregressive nature of LLMs, earlier token information is progressively transferred to later tokens. Skip FFN further enhances this integration by suppressing the attention weights of skipped tokens, forcing the model to focus on key tokens and improving information aggregation efficiency. Empirical results show that, due to the inclusion of summary tokens, this omission does not compromise model performance, underscoring a distinctive advantage of our framework for more efficient inference.
@@ 5 Theoretical analysis of Skip-Vision
Bounded Lipschitz. If the spectral norms of ${W_{1}}$, ${W_{2}}$, ${W_{K}}$, ${W_{Q}}$, and ${W_{V}}$ are bounded by 1, then:
@@ 6.4 Training setup
For LLaVA and LLaVA-HD setting, we use input resolutions of 336 and a dynamic resolution ranging from 336 to 672, respectively, with CLIP ViT-L/336px [87] as the vision encoder and LLaMA3 8B Instruct [29] as the backbone (additional comparisons under LLaVA-1.5-7B training setup are in the Appendix 10.1). For CoS setting, pretraining is conducted at 224 resolution, training only the multi-scale visual resampler with CLIP ViT-L/224px [87]. The model processes 80 visual tokens, including 16 global and 64 local tokens, and scales to 336 and 1296 tokens during SFT. In SFT, the resolution increases to 448, and all model parameters are unfrozen for full optimization.
@@ 7 Conclusion
8 Acknowledgements
This work was supported in part by NSFC (62201342), and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102). This work was supported by Ant Group Research Intern Program.
@@ 10 More experimental results
Following the LLaVA-1.5-7B training setup, we conducted additional comparisons between Skip-Vision and several recent works, as shown in the table 5. MMVet, MMStar and MMBench highlight Skip-Vision’s strength in capturing causal and global information. These benchmarks emphasize high-level reasoning and abstraction, which benefit from Skip-Vision’s ability to preserve essential information flow while reducing redundant computations. By skipping FFN and KV-cache for less informative tokens, the model amplifies signal from key visual cues and enhances causal token interactions. While this comes with a slight trade-off in fine-grained tasks (OCR, Textvqa), it reflects a deliberate balance between perception and reasoning, favoring tasks that rely on semantic integration over detail fidelity.
[…]
Vanilla (576 tokens) 61.9 64.7 58.2 31.1 100%
SparseVLM [131] (64 tokens) 52.7 56.2 51.8 23.3 85.2%
VisionZip [116] (64 tokens) 55.1 60.1 55.5 31.7 93.7%
PDrop [112] (64 tokens) 47.5 58.8 50.6 - -
FasterVLM [128] (58 tokens) 54.9 60.6 55.3 30.1 93.1%
LLaVA-PruMerge [92] (32 tokens) - 60.9 56.0 - -
Skip-Vision ($N_{r}=64,N_{s}=156$) 60.8 65.1 57.4 32.5 100.0%
Table 5: Comparison with more methods